OpenStudy (anonymous):

need help in linear algebra here is the question Select all statements below which are true for all invertible nxn matrices A & B A. A^2 is invertible B. (A+B)(A-B) = A^2-B^2 C. A+A^-1is invertible D. ABA^-1 = B E. (I - A)(I - A) =I -A^2 F. (AB)^-1 = A^-1B^-1 Thank you

OpenStudy (anonymous):

I have trouble with understanding the theory

OpenStudy (anonymous):

If A is invertible, then there exists another matrix A^(-1) (which we call the inverse of A) such that \[A^{-1} A = I \] where I is the identity matrix. If A is invertible, do you think A^2 would be invertible?

OpenStudy (anonymous):

ok i understand

OpenStudy (anonymous):

yes a^2 is invertible

OpenStudy (anonymous):

(A+B)(A-B) = A^2-B^2 this one is no right?

OpenStudy (anonymous):

For the second one, \[(A+B)(A-B) =A^2 -AB+BA-B^2\] so \[(A+B)(A-B) = A^2 - B^2\] if and only if \[AB = BA\] if this is true, we say that A and B commute. Does this appear to have anything to do with invertibility?

OpenStudy (anonymous):

because A^2-B^2 can't satisfy the equation right?

OpenStudy (anonymous):

You are right. The fact that A and B are invertible does not mean that AB = BA.

OpenStudy (anonymous):

ok got it and for the third one

OpenStudy (anonymous):

Well, what do you think?

OpenStudy (anonymous):

A+A^-1--> is not invertible right

OpenStudy (anonymous):

since AA^-1 -> I

OpenStudy (anonymous):

A+A^-1 have nothing to do with invertibility

OpenStudy (anonymous):

If A is invertible, it does not necessarily mean that A + A^(-1) is also invertible. The only such examples I can think of use complex numbers, but that's okay.